8) Show that the graph of +692 +9z²&exy 6x z-12yz-2x - 4y +6z+1=0 is degenerate quadric whizh consists plare. nuestion.
Answer #1

As x²+44² +9 z² +axy - 64 Z-12 4 2 - 2x - ay +63 +1=0 x² + x144 =63-2) +4y² + 9z²1242-44 +6Z+1=0 (²+x (44-62-2) 44 y²49 2² 1212 xº+y149-67-2) 45.99-62-2)] + 4y +93-1212 -44% 9h-x+1242 - 62 + gý - 34 x ² + x 244-62-2) + +62+y=0 (24-37-1)² = 0 (x+2y. 33 X?+4y?+92?) it 4X 7 - 12 YZ - 6x2 = 0 By comparing AX²+2 B xy + (y² + 2DX 2+2EYZ +F 22=0 A=1, B=2,C= 4, D=-3, E =-6,F=9 Co